# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 2.2 Finding a Relevant Direction
Sample and Computational Complexity: From the analysis above, we have that the algorithm terminates in $\mathrm{poly}(1/\epsilon,K)$ iterations and at each iteration we draw $d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$ samples. Hence, we have that the total number of samples is of the order of $d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$. Moreover, the algorithm runs in $\mathrm{poly}(N)$ time as all operations can be implemented in polynomial time. ∎
@@
This section proves Lemma 2.7, one of the supporting lemmas used in the proof of Proposition 2.4. For completeness we restate the lemma. See 2.7 The proof of the lemma immediately follows from Lemma 2.19 and Claim 2.20. First, we establish Lemma 2.19, which provides a bound on the tails of a piecewise affine function. Then, using Claim 2.20, we show that there is indeed a suitable $\gamma$ to which this tail bound can be applied.
Lemma 2.19 (Tail Bounds).
@@
By the properties of addition and scalar multiplication of independent Gaussian random variables, we have that for any $n\in\mathbb{Z}_{+}$ a standard Gaussian random variable $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ can be written as $\mathbf{x}=\frac{1}{n}\sum_{i=1}^{n^{2}}\mathbf{z}^{(i)}$ where $\mathbf{z}^{(i)}$ are drawn i.i.d. from $\mathcal{N}(\mathbf{0},\mathbf{I})$. Now if for some $i\in[n^{2}]$ we have that $s(\mathbf{z}^{(i)})\geq n\gamma$, then there exists a $j\in[K]$ such that $\mathbf{w}^{(j)}\cdot\mathbf{z}^{(i)}+t_{j}\geq n\gamma$ and, furthermore, we have that
$\mathbf{w}^{(j)}\cdot\mathbf{x}+t_{j}$ $=\frac{1}{n}\mathbf{w}^{(j)}\cdot\mathbf{z}^{(i)}+t_{j}+\frac{1}{n}\sum_{k\neq i}\mathbf{w}^{(j)}\cdot\mathbf{z}^{(k)}\geq\gamma+\mathbf{w}^{(j)}\cdot\bigg{(}\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)}\bigg{)}.$
Note that $\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)}$ is distributed according to $\mathcal{N}(\mathbf{0},\frac{n^{2}-1}{n^{2}}\mathbf{I})$; hence, by symmetry we have that $\operatorname*{\mathbf{Pr}}[\mathbf{w}^{(j)}\cdot(\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)})\geq 0]=1/2$. As a result we have that $\operatorname*{\mathbf{Pr}}[s(\mathbf{x})\geq\gamma\mid s(\mathbf{z}^{(i)})\geq n\gamma]\geq 1/2$, for all $i\in[n^{2}]$.
Denote by $p_{n}=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma]$. To prove Lemma 2.19, it suffices to show that $p_{n}\leq 4p_{1}/n^{2}$, for all $n\geq 1$. Note that this holds trivially true for $n=1$. Assume as an inductive hypothesis that $p_{n-1}\leq 4p_{1}/(n-1)^{2}$ for some $n\geq 2$. We will prove that $p_{n}\leq 4p_{1}/n^{2}$ for $n\geq 2$.
@@
where in the first inequality we used the union bound, and in the last inequality we used the independence of the $\mathbf{z}^{(i)}$’s and that $\operatorname*{\mathbf{Pr}}[s(\mathbf{x})\geq\gamma\mid s(\mathbf{z}^{(i)})\geq n\gamma]\geq 1/2$, for all $i\in[n^{2}]$.
@@
Therefore, we have that $p_{n}\leq 4p_{1}/n^{2}$, which by the principle of mathematical induction completes the proof of Lemma 2.19. ∎
In order to apply Lemma 2.19, we need to find a threshold such that the probability that $s$ is greater than this threshold is bounded from above by $1/{48}$. We show the following
@@
Fix $c\in(0,1)$, there exists $\gamma>0$ such that
@@
First notice that the quantity $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ is continuous with respect to $\gamma$ for any $\gamma>0$. Indeed, $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\gamma\leq s(\mathbf{x})\leq\gamma+\epsilon]\leq\sum_{i=1}^{K}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[-t_{i}+\gamma\leq\mathbf{w}^{(i)}\cdot\mathbf{x}\leq-t_{i}+\gamma+\epsilon]$ where each term such that $\mathbf{w}^{(i)}\neq\mathbf{0}$ tends to 00 as $\epsilon$ tends to 00 and each term such that $\mathbf{w}^{(i)}=\mathbf{0}$ is equal to 00 from the fact that $-t_{i}+\gamma>0$. Also note that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ tends to $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]$ as $\gamma$ approaches 00; hence, there exists $\gamma^{\prime}$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]-\epsilon/2<\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma^{\prime}]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]$. Therefore, if $c\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]$ we have that $\gamma^{\prime}$ satisfies the claim, because $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]\geq\epsilon$.
@@
Combining the the above two statements completes the proof of Lemma 2.7.
@@ 3 Learning Multi-Index Models with Agnostic Noise
In this section, we extend the algorithm of Section 2 for a general class of functions that satisfy specific regularity conditions (Definition 1.2), establishing Theorem 1.4. First, we state the formal versions of the conditions and theorem, and then provide pseudocode for the algorithm along with the necessary definitions.
As we will discuss later, we need a continuity condition on the function in order to be able to approximate it over small regions. To this end, we define the Gaussian Surface Area of a multiclass classifier. A bound on this quantity will be crucial for the analysis of our algorithm.
@@
Now we proceed to define the regularity conditions that we will assume in order to get efficient algorithms. Formally, we define a class of well-behaved functions as follows:
@@
Fix $d,K,m\in\mathbb{Z}_{+},\delta,\epsilon,\sigma,\tau\in(0,1)$ and $\mathcal{Y}\subseteq\mathbb{Z}$ of finite cardinality. We say that a class of functions $\mathcal{F}$ from $\mathbb{R}^{d}\mapsto\mathcal{Y}$ is a class of $(m,\epsilon,\delta,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models if, for every $f\in\mathcal{F}$, the following conditions hold:
@@
    Let $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, for any random variable $y$ supported on $\mathcal{Y}$ with $\operatorname*{\mathbf{Pr}}_{\mathbf{x},y}[f(\mathbf{x})\neq y]\leq\epsilon$ and any subspace $V\subseteq\mathbb{R}^{d}$,
@@
      either there exists a function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x},y}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$.
@@
      or with probability at least $\delta$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\mathbf{x}_{0}$ is independent of $\mathbf{x}$, there exists a polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and a label $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x}}[p(\mathbf{x}^{U})]=0$ and $\operatorname*{\mathbf{E}}_{\mathbf{x},y}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>\sigma\|p(\mathbf{x}^{U})\|_{2}$, where $U=(V+W)\cap V^{\perp}$.
We can now state the main theorem of this section.
@@
Fix $d,K,m\in\mathbb{Z}_{+}$, $\zeta,\alpha,\tau,\sigma\in(0,1)$ and $\Gamma\in\mathbb{R}_{+}$. Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\epsilon,\delta\in(0,1)$. Let $\mathcal{F}$ a class of $(m,\zeta,\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2 and let $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$. If $\zeta\geq\mathrm{OPT}+\epsilon$, then, Algorithm 3 draws $N={d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$ i.i.d. samples from $D$, runs in time $\mathrm{poly}(d,N,\lvert\mathcal{Y}\rvert)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon\;.$
Definition 3.2 outlines the crucial properties necessary for our algorithm to produce a satisfactory solution. We first assume that our concept class consists of Multi-Index Models with a bounded Gaussian surface area. This condition bounds the complexity of the decision boundary of the multi-index models or equivalently it bounds how fast the function changes labels. We use this condition in order to show that if we partition our space to sufficiently small cubes then the function should be almost constant. This is a condition that holds for many classes of interest, such as multiclass linear classifiers and intersections of halfspaces.
In addition, we impose a structural condition on the class to ensure progress at each step of the algorithm. Intuitively, if $V$ is the subspace we have currently identified, we require that either it already provides a good approximation Condition 3(a) or we can use moments to uncover further directions in the hidden subspace $W$ Condition 3(b). Specifically, we assume that for any perturbation $y$ of a function $f$ from the class that we are attempting to learn, and for any given subspace $V$, one of the following must hold: (i) there exists a good approximation to $y$ using a function defined over $V$ (Condition 3(a) of Definition 3.2), or (ii) with some probability over orthogonal cylinders to $V$, there exists a low-degree, zero-mean, variance-one polynomial along directions that correlate with the ground-truth subspace $W$, which non-trivially correlates with the level sets of $y$ over that orthogonal cylinder (Condition 3(b) of Definition 3.2). In this manner, extracting these directions and adding them to $V$ enables an improved approximation of the hidden-subspace $W$. Although Condition 3(b) is defined over specific points with a measure of zero, we demonstrate in Lemma 3.5 that, by leveraging our assumptions, there exists a $\delta$ fraction of finite width cylinders that exhibit this property.
@@
Input: Accuracy $\epsilon\in(0,1)$, failure probability $\delta\in(0,1)$, and sample access to a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$.
Output: A hypothesis $h$ such that, with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
@@
    Let $T,p$ be sufficiently large polynomials in $1/\alpha,1/\epsilon,1/\sigma,K,\Gamma,\lvert\mathcal{Y}\rvert,m$, and let $C$ be a sufficiently large universal constant..
@@
      Draw a set $S$ of $N$ i.i.d. samples from $D$.
@@
      $\mathcal{E}_{t}\leftarrow$ Algorithm 4($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S$).
@@
    Return a piecewise constant approximating classifier $h_{\mathcal{S}}$, as defined in Definition 2.3, using $N$ i.i.d. samples from $D$, by finding the most frequent label at any subset of $\mathcal{S}$.
@@
Input: $\epsilon>0$, $\epsilon^{\prime}>0$, $\delta>0$, a subspace $V$ of $\mathbb{R}^{d}$ and a set of samples $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\},\mathbf{x}^{(i)}\in\mathbb{R}^{d},y_{i}\in\mathcal{Y}$.
@@
    For each $S\in\mathcal{S}$ and each $i\in\mathcal{Y}$, find a polynomial $p_{S,i}(\mathbf{x})$ such that
@@
Let $d\in\mathbb{Z}_{+},\epsilon,\epsilon^{\prime}\in(0,1)$ and let $\mathcal{Y}$ be a set of finite cardinality. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$, whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, and let $V\subseteq\mathbb{R}^{d}$ be a $k$-dimensional subspace, with $k\geq 1$. Suppose that for a function $f\in\mathcal{F}$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\epsilon$. Additionally, assume that there exists a function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$. Let $\eta=\epsilon^{\prime 2}/(C\Gamma(f)\lvert\mathcal{Y}\rvert\sqrt{k})$, for $C>0$ a sufficiently large universal constant, and $\mathcal{S}$ be an $\eta$-piecewise approximating partition with respect to $V$, as defined in Definition 2.2. Then, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]\leq\tau+\epsilon+\epsilon^{\prime}$, where $h_{\mathcal{S}}$ is an $\eta$-piecewise constant approximation function as defined in Definition 2.3.

### v2

@@ 2.2 Finding a Relevant Direction
Sample and Computational Complexity: From the analysis above, we have that the algorithm terminates in $T=\mathrm{poly}(1/\epsilon,K)$ iterations and at each iteration we draw $d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$ samples. Hence, we have that the total number of samples is of the order of $d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$. Moreover, the algorithm runs in $\mathrm{poly}(N)$ time as all the computational operations can be implemented in polynomial time. ∎
@@
This section proves Lemma 2.7, one of the supporting technical lemmas used in the proof of Proposition 2.4. For completeness we restate the lemma below. See 2.7
Proof.
The proof of the lemma immediately follows from Claim 2.19 and Claim 2.20. First, we establish Claim 2.19, which provides a bound on the tails of a piecewise affine function. Then, using Claim 2.20, we show that there is indeed a suitable $\gamma$ to which this tail bound can be applied.
Claim 2.19 (Tail Bounds).
@@
By the properties of addition and scalar multiplication of independent Gaussian random variables, we have that for any $n\in\mathbb{Z}_{+}$ a standard Gaussian random variable $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ can be written as $\mathbf{x}=\frac{1}{n}\sum_{i=1}^{n^{2}}\mathbf{z}^{(i)}$ where $\mathbf{z}^{(i)}$ are drawn i.i.d. from $\mathcal{N}(\mathbf{0},\mathbf{I})$. Now if for some $i\in[n^{2}]$ we have that $s(\mathbf{z}^{(i)})\geq n\gamma$, then there exists $j\in[K]$ such that $\mathbf{w}^{(j)}\cdot\mathbf{z}^{(i)}+t_{j}\geq n\gamma$ (by definition of $s(\mathbf{x})$) and, furthermore, we have that
$\mathbf{w}^{(j)}\cdot\mathbf{x}+t_{j}$ $=\frac{1}{n}\mathbf{w}^{(j)}\cdot\mathbf{z}^{(i)}+t_{j}+\frac{1}{n}\sum_{k\neq i,k\in[n^{2}]}\mathbf{w}^{(j)}\cdot\mathbf{z}^{(k)}\geq\gamma+\mathbf{w}^{(j)}\cdot\bigg{(}\frac{1}{n}\sum_{k\neq i,k\in[n^{2}]}\mathbf{z}^{(k)}\bigg{)}.$
Note that $\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)}$ is distributed according to $\mathcal{N}(\mathbf{0},\frac{n^{2}-1}{n^{2}}\mathbf{I})$. Hence, by symmetry we have that $\operatorname*{\mathbf{Pr}}[\mathbf{w}^{(j)}\cdot(\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)})\geq 0]=1/2$. As a result we have that $\operatorname*{\mathbf{Pr}}[s(\mathbf{x})\geq\gamma\mid s(\mathbf{z}^{(i)})\geq n\gamma]\geq 1/2$, for all $i\in[n^{2}]$.
Denote by $p_{n}=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma]$. To prove Claim 2.19, it suffices to show that $p_{n}\leq 4p_{1}/n^{2}$, for all $n\geq 1$. We will prove it by induction. Note that for $n=1$ the hypothesis holds trivially. Assume as an inductive hypothesis that $p_{n-1}\leq 4p_{1}/(n-1)^{2}$ for some $n\geq 2$. We will prove that $p_{n}\leq 4p_{1}/n^{2}$ for $n\geq 2$.
@@
where in the first inequality we used the union bound, and in the second inequality we used the independence of the $\mathbf{z}^{(i)}$’s and that $\operatorname*{\mathbf{Pr}}[s(\mathbf{x})\geq\gamma\mid s(\mathbf{z}^{(i)})\geq n\gamma]\geq 1/2$, for all $i\in[n^{2}]$.
@@
Therefore, we have that $p_{n}\leq 4p_{1}/n^{2}$, which completes the proof of Claim 2.19. ∎
To apply Claim 2.19, we must choose a threshold $\gamma$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(0,I)}[s(\mathbf{x})\geq\gamma]\leq 1/48.$ We now show the following result.
@@
For any sufficiently small $c\in(0,1)$, there exists $\gamma>0$ such that
@@
Notice that the quantity $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ is continuous with respect to $\gamma$ for any $\gamma>0$. Indeed, $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\gamma\leq s(\mathbf{x})\leq\gamma+\epsilon]\leq\sum_{i=1}^{K}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[-t_{i}+\gamma\leq\mathbf{w}^{(i)}\cdot\mathbf{x}\leq-t_{i}+\gamma+\epsilon]$ where each term such that $\mathbf{w}^{(i)}\neq\mathbf{0}$ tends to 00 as $\epsilon$ tends to 00 and each term such that $\mathbf{w}^{(i)}=\mathbf{0}$ is equal to 00 from the fact that $-t_{i}+\gamma>0$. Also note that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ tends to $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]$ as $\gamma$ approaches 00; hence, there exists $\gamma^{\prime}$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]-\epsilon/2<\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma^{\prime}]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]$. Therefore, if $c\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]$ we have that $\gamma^{\prime}$ satisfies the claim, because $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]\geq\epsilon$.
@@
Combining the the above two statements completes the proof of Lemma 2.7. ∎
@@ 3 Learning Multi-Index Models with Agnostic Noise
In this section, we extend the algorithm presented in Section 2 for a general class of MIMs that satisfy specific regularity conditions (Definition 1.2), establishing Theorem 1.5. We begin by stating the formal versions of the regularity conditions and the theorem. Following this, we provide the pseudocode for the extended algorithm along with the necessary definitions.
As we will discuss later, we need a continuity condition on the target function in order to be able to accurately approximate it over small regions. To this end, we define the Gaussian Surface Area of a multiclass classifier. A bound on this quantity will be crucial for the analysis of our algorithm.
@@
Now we proceed to define the regularity conditions that we assume in order to obtain efficient algorithms.
@@
Let $d,K,m\in\mathbb{Z}_{+},\alpha,\zeta,\sigma,\tau\in(0,1)$ and $\mathcal{Y}\subseteq\mathbb{Z}$ of finite cardinality. We say that a class of functions $\mathcal{F}$ from $\mathbb{R}^{d}\mapsto\mathcal{Y}$ is a family of concept classes of $(m,\zeta,\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models if, for every $f\in\mathcal{F}$, the following conditions hold:
@@
    Let $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. For any random variable $y$ supported on $\mathcal{Y}$ with $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]\leq\zeta$ and any subspace $V\subseteq\mathbb{R}^{d}$,
@@
      either there exists a function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x},y}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$,
@@
      or with probability at least $\alpha$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\mathbf{x}_{0}$ is independent of $\mathbf{x}$, there exists a polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and a label $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}[p(\mathbf{x}^{U})]=0$ and $\operatorname*{\mathbf{E}}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>\sigma\|p(\mathbf{x}^{U})\|_{2}$, where $U=(V+W)\cap V^{\perp}$.
We can now state the formal version of Theorem 1.5.
@@
Fix $d,K,m\in\mathbb{Z}_{+}$, $\zeta,\alpha,\tau,\sigma\in(0,1)$ and $\Gamma\in\mathbb{R}_{+}$. Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\epsilon,\delta\in(0,1)$. Let $\mathcal{F}$ a class of $(m,\zeta,\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2 and let $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$. If $\zeta\geq\mathrm{OPT}+\epsilon$, then, Algorithm 3 draws $N={d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$ i.i.d. samples from $D$, runs in time $\mathrm{poly}(N)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon\;.$
Definition 3.2 outlines the crucial properties necessary for our algorithm to produce a satisfactory solution. We first assume that our concept class consists of Multi-Index Models with finite Gaussian surface area. This condition bounds the complexity of the decision boundary of the multi-index models. We leverage this condition in order to show that if we partition our space to sufficiently small cubes then the function should be almost constant within each cube. This is a condition that holds for many classes of interest, such as multiclass linear classifiers and intersections of halfspaces.
In addition, we impose a structural condition on the target class to guarantee progress at each step of the algorithm. Intuitively, if $V$ is the subspace we have currently identified, we require that either it already provides a good approximation (Condition 3(a)) or we can use moments to uncover further directions in the hidden subspace $W$ Condition 3(b). Specifically, we assume that for any perturbation $y$ of a function $f$ from the class that we are attempting to learn, and for any given subspace $V$, one of the following must hold: (i) there exists a good approximation to $y$ using a function defined over $V$ (Condition 3(a)), or (ii) with some probability over orthogonal cylinders to $V$, there exists a low-degree, zero-mean, variance-one polynomial along directions that correlate with the ground-truth subspace $W$, which non-trivially correlates with the level sets of $y$ over that orthogonal cylinder (Condition 3(b). In this manner, extracting these directions and adding them to $V$ enables an improved approximation of the hidden-subspace $W$. Although Condition 3(b) is defined over specific points with a measure of zero, we demonstrate in Lemma 3.5 that, by leveraging our assumptions, there exists a $\alpha$ fraction of finite width cylinders that exhibit this property.
@@
Input: Accuracy $\epsilon\in(0,1)$, failure probability $\delta\in(0,1)$, and sample access to a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
Output: A hypothesis $h$ such that, with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$, where $\mathrm{OPT}\coloneqq\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
@@
    Let $T,p$ be sufficiently large polynomials in $1/\alpha,1/\epsilon,1/\sigma,K,\Gamma,\lvert\mathcal{Y}\rvert,m$, and let $C$ be a sufficiently large universal constant.
@@
      Draw a set $S_{t}$ of $N$ i.i.d. samples from $D$.
@@
      $\mathcal{E}_{t}\leftarrow$ Algorithm 4($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S_{t}$).
@@
    Draw $N$ i.i.d. samples from $D$ and construct the piecewise constant classifier $h_{\mathcal{S}}$ as follows: For each $S\in\mathcal{S}$, assign a label that appears most frequently among the samples falling in $S$. Return $h_{\mathcal{S}}$.
@@
Input: $\epsilon>0$, $\epsilon^{\prime}>0$, $\delta>0$, a subspace $V$ of $\mathbb{R}^{d}$ and a set of samples $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\}$ from a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@
    For each $S\in\mathcal{S}$ and each $i\in\mathcal{Y}$, compute a polynomial $p_{S,i}(\mathbf{x})$ such that
@@
Let $d,k\in\mathbb{Z}_{+},\epsilon,\epsilon^{\prime}\in(0,1)$ and let $\mathcal{Y}$ be a set of finite cardinality. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$, whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, and let $V\subseteq\mathbb{R}^{d}$ be a $k$-dimensional subspace, with $k\geq 1$. Suppose that for a function $f\in\mathcal{F}$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\epsilon$. Additionally, assume that there exists a function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$. Let $\eta=\epsilon^{\prime 2}/(C\Gamma(f)\lvert\mathcal{Y}\rvert\sqrt{k})$, for $C>0$ a sufficiently large universal constant, and $\mathcal{S}$ be an $\eta$-piecewise approximating partition with respect to $V$ (see Definition 2.2). Then, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]\leq\tau+\epsilon+\epsilon^{\prime}$, where $h_{\mathcal{S}}$ is an $\eta$-piecewise constant approximation function (see Definition 2.3).